Last year, Joan purchased a $1,000 face value corporate bond with an 7% annual coupon rate and a 25-year maturity. At the time of the purchase, it had an expected yield to maturity of 12.49%. If Joan sold the bond today for $1,018.1, what rate of return would she have earned for the past year? Round your answer to two decimal places.
Answer #1

Answer:

Rate of return earned in the past year=8.81%

face value $                 1,000.00
annual coupon rate 7%
maturity 25
sale price $                 1,018.10
Formula:
return on investment=(sale price+receipts-purchase price)/purchase price x100
=(1018.10+(face valuexcoupon rate)-1000)/1000x100
=(18.1+(1000*7%))/1000*100
=(18.1+70)/1000x100
=88.1/1000x100
8.81%